Corrigendum

Cohomology and Euler characteristics of Coxeter groups

Toshiyuki AKITA^{∗}

Department of Applied Mathematics, Fukuoka University, Fukuoka 814-0180, Japan

The inequality e(W) ≥ γ(F) in Theorem 7 was stated incorrectly. The correct inequality is

e(W) ≤ γ(F).

Similarly, the inequality e(W) ≥ γ(F(W)) at the bottom of p. 3 was incorrect. It should be

e(W) ≤ γ(F(W)).

∗This paper was originally published in Science Bulletin of Josai University, Special Issue No. 2 (1997), 3–16.